# Laser-FLASH: radiobiology at high dose, ultra-high dose-rate, single pulse laser-driven proton source

### v1

@@ Abstract
biological samples. A transport line composed by two permanent magnet quadrupoles
@@
protocol[1], are presented for in vitro and
in vivo irradiation are presented. A reduction of radiation-induced oxidative
@@ Introduction
The recent discovery of the FLASH effect by Favaudon et al.[7] renewed the interest around the role of dose-rate and temporal modality of dose deposition in defining the biological and the physiological effect of ionizing radiation. Some advantages demonstrated for in-vivo and human irradiation (see Wilson et al.[8] and references therein for a consistent review) suggest a possible improvement of the therapeutic window at high dose rate, which motivates the exploration of non-conventional temporal dose deposition modalities. Laser-driven particle sources produce short and bright particle bunches, as each laser pulse independently extracts and accelerates a short and bright packet. Owing to the ultra-fast nature of laser pulses at ultra-relativistic intensity, laser-accelerated protons have a duration, at the source, between few picoseconds and few nanoseconds[9, 10].
@@
within a single laser-driven proton pulse. A set of permanent magnet quadrupoles is used
to transport and shape the protons to an irradiation plane in air, far from the
@@
The experiment was performed in two separate campaigns at the LULI laboratory (Palaiseau, France) on the pico2000 laser system. The picosecond beam (duration , energy per pulse, ) is focused by a off-axis parabola (PFL , d=).
@@
by two permanent magnet quadrupoles and a variable-thickness scattering filter.
@@
The role of the transport line is to control the particle density and, consequently, the deposited dose at the target plane, far from the laser-plasma interaction point. During the design step the deposited dose at the irradiation plane is simulated by Monte Carlo methods (Geant4 toolkit, see Methods section) for different beam line configurations. The source is modeled as being purely thermal with a high energy cutoff and no angular dependence. This assumption is based on the small effective aperture of the first quadrupole, which acts as a filter on the broad divergent source. The largest accepted half-width angle () is, according to measurements in Figure 2, smaller than the source divergence at the cutoff energy. For this reason we decided to drop the angular dependence in the source modelling; the useful spectral charge is then represented analytically as:
@@ Discussion
organ-at-risk. The Bragg pick dose deposition profile allows the dose to be better
@@ Methods
The complete beamline is simulated using the Monte Carlo toolkit Geant4, with quadrupoles
represented
by field maps. The possibility of simulating magnetic optics with TNSA
@@
in equation 1, where lower and higher cutoffs matching experimental
@@
The cell containers used for irradiation were lumox©dish 35 (SARSTEDT) exhibiting a thick lumox©bottom face. Depending on beam transverse profile and position observed on radiochromic films (ref) a circular area was delimited on the internal face of the lumox©membrane, where cells were seeded. Cells were let grow overnight in of medium. To generate dose-response survival curves U87-MG cell lines were subjected to doses varying from with single bunch of laser driven protons. Survival curves resulting from irradiation with CAP or fast-fractionation LDP were retrieved from [4]. After exposure to ionizing radiations, cells were incubated for in standard conditions. Cells were harvested with Accutase (Merck), dispatched into different wells of -well plates (TPP) in of medium and grown for five generations corresponding to for U87-MG cell line. Cells were harvested with Accutase which was then inactivated using an equal volume of PBS (ThermoFisher Scientific) supplemented with fetal calf serum. The final volume was adjusted to with PBS and of each well were dispatched into a non-sterile U-bottom -well plate (TPP). In each well, of a propidium iodide solution (Sigma, 100 in 1X PBS) were added just before flow cytometry counting. Cell acquisition and data analysis were performed using Guava©and GuavaSoft (Merck) and then GraphPad Prism software.

### v2

@@ Abstract
biological samples. A transport line composed by two permanent-magnet quadrupoles
@@
protocol[1], are presented for in vitro and in vivo irradiation. A reduction of radiation-induced oxidative
@@ Introduction
The recent discovery of the FLASH effect by Favaudon et al.[7] renewed the interest around the role of dose-rate and temporal modality of dose deposition in defining the biological and the physiological effect of ionizing radiation. Some advantages demonstrated for in-vivo and human irradiation (see Wilson et al.[8] and references therein for a consistent review) suggest a possible improvement of the therapeutic window at high dose rate, which motivates the exploration of non-conventional temporal dose deposition modalities. Laser-driven particle sources produce short and bright particle bunches, as each laser pulse independently extracts and accelerates a short and bright packet. Owing to the ultra-fast nature of laser pulses at relativistic intensity, laser-accelerated protons have a duration, at the source, between few picoseconds and few nanoseconds[9, 10].
@@
within a single laser-driven proton pulse. A set of permanent-magnet quadrupoles is used
to transport and shape the protons to an irradiation area in air, far from the
@@
The experiment was performed in two separate campaigns at the LULI laboratory (École Polytechnique, Palaiseau, France) on the pico2000 laser system. The picosecond beam (duration , energy per pulse, ) is focused by a off-axis parabola (PFL , d=).
@@
by two permanent-magnet quadrupoles and a variable-thickness scattering filter.
@@
The role of the transport line is to control the particle density and, consequently, the deposited dose at the target plane, far from the laser-plasma interaction point. During the design step the deposited dose at the irradiation plane is simulated by Monte Carlo methods (Geant4 toolkit) for different beam line configurations. The source is modeled as being purely thermal with a high energy cutoff and no angular dependence. This assumption is based on the small effective aperture of the first quadrupole, which acts as a filter on the broad divergent source. The largest accepted half-width angle () is, according to measurements in Figure 2, smaller than the source divergence at the cutoff energy. For this reason we decided to drop the angular dependence in the source modelling; the useful spectral charge is then represented analytically as:
@@ Discussion
organ-at-risk. The Bragg peak dose deposition profile allows the dose to be better
@@ Methods
The proton beamline is simulated using the Monte Carlo toolkit Geant4, where quadrupoles
are represented
by measured field maps. The possibility of simulating magnetic optics with TNSA
@@
in equation (1), where lower and higher cutoffs matching experimental
@@
The cell containers used for irradiation were lumox® dish 35 (SARSTEDT) exhibiting a thick lumox® bottom face. Depending on beam transverse profile and position observed on radiochromic films (ref) a circular area was delimited on the internal face of the lumox® membrane, where cells were seeded. Cells were let grow overnight in of medium. To generate dose-response survival curves U87-MG cell lines were subjected to doses varying from with single bunch of laser driven protons. Survival curves resulting from irradiation with CAP or fast-fractionation LDP were retrieved from [4]. After exposure to ionizing radiations, cells were incubated for in standard conditions. Cells were harvested with Accutase (Merck), dispatched into different wells of -well plates (TPP) in of medium and grown for five generations corresponding to for U87-MG cell line. Cells were harvested with Accutase which was then inactivated using an equal volume of PBS (ThermoFisher Scientific) supplemented with fetal calf serum. The final volume was adjusted to with PBS and of each well were dispatched into a non-sterile U-bottom -well plate (TPP). In each well, of a propidium iodide solution (Sigma, 100 in 1X PBS) were added just before flow cytometry counting. Cell acquisition and data analysis were performed using Guava® and GuavaSoft (Merck) and then GraphPad Prism software.